# Best Agent Identification for General Game Playing

### v1

@@ 5.4 Discussion
Overall, the improvements in best agent identification accuracy and efficiency offered by our proposed Optimistic-WS algorithm are clear across both domains, being able to to consistently outperform all other techniques for any number of rounds tested. The benefits of our Optimistic-WS algorithm appear most prominent in the early-mid round period, where it quickly achieved a very low simple regret compared to other techniques. As more rounds were performed we see that this gap begins to shrink, although the overall percentage improvement remains relatively stable. Given an infinite number of rounds, all techniques would (in theory) be able to obtain a perfect average simple regret score of zero, although such an assumption is clearly not feasible in practice. Demonstrating our Optimistic-WS algorithm’s improved performance over a wide range of possible rounds emphasises its practical application to estimating agent performance in a realistic experimental setting.
6 Conclusion
@@
In this paper we have presented our novel Optimistic-WS algorithm for best arm identification in multi-armed bandits. This approach uses the Wilson score interval to estimate the regret change potential of each arm, and has been designed specifically for the task of identifying the best agents in a general game playing domain. Experiments conducted on two of the most popular general game systems (GVGAI and Ludii) reveal that our proposed approach significantly outperforms all previous state-of-the-art algorithms. We hope that this approach will be applied to enhance future agent evaluation procedures across a wide range of general game systems.
6.1 Future Work
@@
There were several aspects that we did not consider for our presented Optimistic-WS algorithm that could be explored further. In our experiments, all trials were assumed to take an equal length of time to complete. However, this is often not the case, with certain games likely taking much longer to complete than others. Skilled agents may also be able to win easy games quicker than less skilled agents, or conversely take longer to lose at harder games. While we did not explore these aspects within this paper, it should be possible to create a “cost aware” version of our algorithm that considers the expected computational expense of performing a trial for a specific game-agent pair (based on prior experience). We also did not consider any domain specific measures that could be used to identify performance correlations between games/agents. If we are able to estimate that two (or more) games/agents have similar profiles, then results obtained from one trial could potentially be applied proportionately to several other related games/agents. Domain agnostic measures of game/agent distance could also be determined based on the similarity of prior results, although these may not always prove reliable.
@@
Another suitable avenue for future work would be to identify additional multi-task, multi-agent domains for carrying out further evaluations. Our approach is general enough to work on datasets for any such domains, providing the requirement on reward limits being bounded between 0 and 1 is satisfied. While we selected the Wilson score confidence interval specifically for the domain of general game playing, it is also possible that other confidence intervals would be better suited for alternative domains with different reward distributions. Further evaluations will therefore be needed to accurately estimate the wider applicability of our approach.
@@
A GVGAI (all levels) additional hyperparameter results
@@
In this appendix we provide additional results for other algorithm hyperparameter values not shown in Figure 1
@@
Figure 4: Average simple regret for each GapE hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 5: Average simple regret for each GapE-V hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 6: Average simple regret for each UCB-E hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 7: Average simple regret for each Optimistic-WS hyperparameter value when applied to the GVGAI (all levels) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
B GVGAI (level 1) additional hyperparameter results
In this appendix we provide additional results for other algorithm hyperparameter values not shown in Figure 2
Figure 8: Average simple regret for each GapE hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
Figure 9: Average simple regret for each GapE-V hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@ B GVGAI (level 1) additional hyperparameter results
Figure 10: Average simple regret for each UCB-E hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 11: Average simple regret for each Optimistic-WS hyperparameter when applied to the GVGAI (level 1) game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
C Ludii additional hyperparameter results
@@
In this appendix we provide additional results for other algorithm hyperparameter values not shown in Figure 3
@@
Figure 12: Average simple regret for each GapE hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 13: Average simple regret for each GapE-V hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 14: Average simple regret for each UCB-E hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.
@@
Figure 15: Average simple regret for each Optimistic-WS hyperparameter when applied to the Ludii game-agent results dataset. The shaded uncertainty region indicates plus/minus one standard deviation.

### v2

@@ VI-F Discussion
The benefits of our RCP algorithm also appear to be most prominent in the early-mid round period, where it quickly achieved a much lower simple regret compared to other techniques. As more rounds are performed we see that this gap begins to shrink, although the overall percentage improvement remains relatively stable. Given an sufficiently large number of rounds, even random arm selection would (in theory) be able to obtain a perfect average simple regret score of zero, although such an assumption is clearly not feasible in practice. Demonstrating our RCP algorithm’s improved performance over a wide range of possible rounds emphasises its practical application to estimating agent performance in a realistic experimental setting.
VII Conclusion
@@
In this paper we have presented our novel RCP algorithm for best arm identification in multi-armed bandits. This approach uses a provided confidence interval to estimate the regret change potential of each arm, and has been designed specifically for the task of identifying the best agents in a general game playing domain. Experiments conducted on two of the most popular general game systems (GVGAI and Ludii) reveal that our proposed approach significantly outperforms all previous state-of-the-art algorithms. We hope that this approach will be applied to enhance future agent evaluation procedures across a wide range of general game systems.
VII-A Future Work
@@
There are several aspects that we did not consider for our presented RCP algorithm that could be explored further. In our experiments, all trials were assumed to take an equal length of time to complete. However, this is often not the case, with certain games likely taking much longer to complete than others. Skilled agents may also be able to win easy games quicker than less skilled agents, or conversely take longer to lose at harder games. While we did not explore these aspects within this paper, it should be possible to create a “cost aware” version of our algorithm that considers the expected computational expense of performing a trial for a specific game-agent pair (based on prior experience). We also did not consider any domain specific measures that could be used to identify performance correlations between games/agents. If we are able to estimate that two (or more) games/agents have similar profiles, then results obtained from one trial could potentially be applied proportionately to several other related games/agents. Domain agnostic measures of game/agent distance could also be determined based on the similarity of prior results, although these may not always prove reliable.
One limitation with some of our proposed confidence interval estimates (specifically Wilson Score, Jeffreys, Clopper–Pearson and Agresti–Coull) is that they do not consider the total number of samples (i.e., rounds) that have been performed. This makes them fast to compute, as only the confidence bounds of the last arm pulled needs to be updated, but also means that the confidence bounds of an arm only change if it is pulled. This could lead to a situation where, through very bad luck, the lower confidence bound of a non-best arm exceeds the upper confidence bound of the best arm. While this is very unlikely to happen, particularly for very small alpha values, it is still theoretically possible. One way to address this problem would be to use an adaptive alpha value that decreases as the current round number increases (similar to that of PAC). A correctly tuned adaptation rate could also lead to improved performance, by starting off with a high alpha value that finds good (although potentially not optimal) solutions quickly, and then steadily decreasing this alpha value to increase the amount of exploration (i.e., reverse annealing).
@@
Another suitable avenue for future work would be to identify additional multi-task, multi-agent domains for carrying out further evaluations. Our approach is general enough to work on datasets for any such domains, providing the requirement on reward limits being bounded between 0 and 1 is satisfied. While we selected binomial confidence interval estimates specifically for the domain of general game playing, it is also possible that other confidence intervals would be better suited for alternative domains with different reward distributions. Further evaluations will therefore be needed to accurately estimate the wider applicability of our approach.
@@
Appendix A GVGAI (all levels) additional hyperparameter results
@@
In this appendix we provide additional results for all algorithm hyperparameter values on the GVGAI (all levels) dataset.
@@
Figure 13: Average simple regret for each GapE exploration value when applied to the GVGAI (all levels) game-agent results dataset.
@@
Figure 14: Average simple regret for each GapE-V exploration value when applied to the GVGAI (all levels) game-agent results dataset.
@@
Figure 15: Average simple regret for each UCB-E exploration value when applied to the GVGAI (all levels) game-agent results dataset.
@@
Figure 16: Average simple regret for each RCP confidence interval ($\alpha$ = 0.1) when applied to the GVGAI (all levels) game-agent results dataset.
@@
Figure 17: Average simple regret for each RCP confidence interval ($\alpha$ = 0.05) when applied to the GVGAI (all levels) game-agent results dataset.
@@
Figure 18: Average simple regret for each RCP confidence interval ($\alpha$ = 0.01) when applied to the GVGAI (all levels) game-agent results dataset.
Appendix B GVGAI (level 1) additional hyperparameter results
@@
In this appendix we provide additional results for all algorithm hyperparameter values on the GVGAI (level 1) dataset.
@@
Figure 19: Average simple regret for each GapE hyperparameter when applied to the GVGAI (level 1) game-agent results dataset.
@@
Figure 20: Average simple regret for each GapE-V hyperparameter when applied to the GVGAI (level 1) game-agent results dataset.
@@
Figure 21: Average simple regret for each UCB-E hyperparameter when applied to the GVGAI (level 1) game-agent results dataset.
@@
Figure 22: Average simple regret for each RCP confidence interval ($\alpha$ = 0.1) when applied to the GVGAI (level 1) game-agent results dataset.
Figure 23: Average simple regret for each RCP confidence interval ($\alpha$ = 0.05) when applied to the GVGAI (level 1) game-agent results dataset.
Figure 24: Average simple regret for each RCP confidence interval ($\alpha$ = 0.01) when applied to the GVGAI (level 1) game-agent results dataset.
Appendix C Ludii additional hyperparameter results
In this appendix we provide additional results for all algorithm hyperparameter values on the Ludii dataset.
Figure 25: Average simple regret for each GapE hyperparameter when applied to the Ludii game-agent results dataset.
Figure 26: Average simple regret for each GapE-V hyperparameter when applied to the Ludii game-agent results dataset.
Figure 27: Average simple regret for each UCB-E hyperparameter when applied to the Ludii game-agent results dataset.
Figure 28: Average simple regret for each RCP confidence interval ($\alpha$ = 0.1) when applied to the Ludii game-agent results dataset.
Figure 29: Average simple regret for each RCP confidence interval ($\alpha$ = 0.05) when applied to the Ludii game-agent results dataset.
Figure 30: Average simple regret for each RCP confidence interval ($\alpha$ = 0.01) when applied to the Ludii game-agent results dataset.
Appendix D Probably Approximately Correct Result
We present a standard fixed-confidence guarantee for our the RCP algorithm when employing the previously described PAC confidence interval. This guarantee relies on (i) confidence intervals that hold uniformly over time and (ii) a separation-based stopping rule.
For each bandit $m\in\{1,\dots,M\}$ and arm $k\in\{1,\dots,K\}$, rewards are i.i.d. random variables in $[0,1]$ with mean $\mu_{m,k}$. Recall that $T_{m,k}(t)$ denotes the number of observations collected from pair $(m,k)$ up to time $t$, and let $\hat{\mu}_{m,k}(t)$ be the corresponding sample mean. For a user-defined confidence parameter $\delta\in(0,1)$, define
$\delta_{m,k,t}:=\frac{6\delta}{\pi^{2}MKt^{2}},$ (7)
the confidence radius
$\beta_{m,k}(t):=\sqrt{\frac{\log\!\bigl(2/\delta_{m,k,t}\bigr)}{2\max\{1,T_{m,k}(t)\}}},$ (8)
and the confidence interval $L_{m,k}(t),U_{m,k}(t)]$ with
$L_{m,k}(t):=\max\{0,\hat{\mu}_{m,k}(t)-\beta_{m,k}(t)\},$ (9)
$U_{m,k}(t):=\min\{1,\hat{\mu}_{m,k}(t)+\beta_{m,k}(t)\}.$ (10)
Let $\hat{k}_{m}(t)\in\arg\max_{k}\hat{\mu}_{m,k}(t)$ be the estimated best arm on bandit $m$ at time $t$. We use a stopping rule where $\tau$ is the first time $t$ such that, for every $m$,
$L_{m,\hat{k}_{m}(t)}(t)\ \geq\ \max_{k\neq\hat{k}_{m}(t)}U_{m,k}(t).$ (11)
Upon stopping, we make final recommendations of $J_{m}:=\hat{k}_{m}(\tau)$ for each $m$.
Define the event that all confidence intervals contain the true means simultaneously for all times:
$\mathcal{E}:=\bigcap_{t\geq 1}\ \bigcap_{m=1}^{M}\ \bigcap_{k=1}^{K}\{\mu_{m,k}\in[L_{m,k}(t),U_{m,k}(t)]\}.$ (12)
Lemma 1
With rewards bounded within $[0,1]$, we have $\Pr(\mathcal{E})\geq 1-\delta$.
Proof:
Fix any triple $(m,k,t)$. The random variable $\hat{\mu}_{m,k}(t)$ is the average of $n$ i.i.d. samples in $[0,1]$ with mean $\mu_{m,k}$. By Hoeffding’s inequality, for any $n\geq 1$,
$\Pr\!\left(|\hat{\mu}_{m,k}(t)-\mu_{m,k}|>\sqrt{\frac{\log(2/\delta_{m,k,t})}{2n}}\ \Big|\ T_{m,k}(t)=n\right)$
$\leq\delta_{m,k,t}.$ (13)
The definition (8) uses $\max\{1,T_{m,k}(t)\}$ so the same bound holds unconditionally, implying
$\Pr\big(\mu_{m,k}\notin[L_{m,k}(t),U_{m,k}(t)]\big)\leq\delta_{m,k,t}.$
Then applying a union bound over all $m,k,t$:
$\Pr(\neg\mathcal{E})\leq\sum_{t=1}^{\infty}\sum_{m=1}^{M}\sum_{k=1}^{K}\delta_{m,k,t}=\sum_{t=1}^{\infty}MK\cdot\frac{6\delta}{\pi^{2}MKt^{2}}$
$=\frac{6\delta}{\pi^{2}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}=\frac{6\delta}{\pi^{2}}\cdot\frac{\pi^{2}}{6}=\delta.$ (14)
Assuming that the stopping rule has been reached after some amount of time, we have that for all bandits $m$,
$\mu_{m,\hat{k}}\ \geq\ L_{m,\hat{k}}(t)\ \geq\ \max_{k\neq\hat{k}}U_{m,k}(t)\ \geq\ \max_{k\neq\hat{k}}\mu_{m,k}.$
Hence $\mu_{m,\hat{k}}=\max_{k}\mu_{m,k}$, that is, $\hat{k}\in\arg\max_{k}\mu_{m,k}$. Therefore the probability of making at least one incorrect recommendation satisfies
$\Pr\big(\exists m:\ J_{m}\notin\arg\max_{k}\mu_{m,k}\big)\leq\delta.$
Acknowledgment
This work was supported by the European Cooperation in Science and Technology (COST) Action #CA22145.